OpenStudy (anonymous):

anybody help in lab2 section of the mitx

OpenStudy (anonymous):

hope u will get

OpenStudy (anonymous):

The goal is to design a mixer circuit with characteristics of Vout=1/2*V1+1/6*V2 You might've started to design your mixer with two resistors only as the example suggests. But working through the math, soon you'll realize that the equations return no non-zero value for the resistor components. Thus you have to change the design. The next simplest design will be to add a resistor (R3) that connects the node Vout to ground. See the schematic below: Since we're going to use linear elements in this circuit (merely linear resistors), therefore superposition will hold and thus one can look at the effect of each source V1 and V2 one at the time: • V1 effect: Vout 1 = V1 * (R2||R3)/(R2||R3+R1) • V2 effect: Vout 2 = V2 * (R1||R3)/(R1||R3+R2) Vout = Vout 1 + Vout 2 = Vout = V1 * (R2||R3)/(R2||R3+R1) + V2 * (R1||R3)/(R1||R3+R2) = 1/2*V1+1/6*V2 Therefore: • (R2||R3)/(R2||R3+R1) = 1/2 (1) • (R1||R3)/(R1||R3+R2) = 1/6 (2) So we have to solve for the resistors given these two equations. (Of course, 2 equations, 3 unknowns will yield one of them to be free to choose, i.e. one will get other two resistors in terms of the third.) • Equation (1) results in R1 = R2||R3 • And equation (2) results in R2 = 5 * R1||R3 Spanning out these two equations one will get: • 1/R1=1/R2+1/R3 (3) • 1/R1+1/R3=5/R2 (4) Subtracting these two equations will yield R2 = 2 * R3 And putting this back to equation (3), will result in R1 = 2/3 * R3 So now we have R2 and R1 in terms of R3 with the following ratios: • R2 = 2 * R3 • R1 = 2/3 * R3 Since the design hadn't mentioned anything about the resistances, one can use a simple value of R3= 3Ω and find the rest accordingly: • R1= 2Ω • R2= 6Ω • R3= 3Ω With the resistor values mentioned above, the transient response is the following, which exactly meets the required specs of Vout. Also the schematic of the final circuit is given for the reference. (Note: Any other values that preserve these ratios between resistors will work as well. For instance for a lower current case one might choose R1= 2kΩ, R2= 6kΩ and R3= 3kΩ